OpenStudy (anonymous):

A uniform solid cylinder of diameter 14cm rolls with out slipping with angular velocity 25 rad/s . The cylinder suddenly contacts a plane inclined at an angle theta. The Value of theta which Brings the cylinder immediately to rest is ?

OpenStudy (anonymous):



OpenStudy (anonymous):

won't 90 degrees do?

OpenStudy (anonymous):

oh wait.. that ll just bounce it back i guess

OpenStudy (anonymous):

never mind!

OpenStudy (shubhamsrg):

 Linear velocity before = wr = 25*0.07 = 1.75 let it remain in contact with plane for time dt, then, (Nsinx)dt = m(1.75) also, it stops rolling, then Tdt = I(25) where T=torque => (Ncosx)(0.07) dt= (1/2)m(0.07)^2 * 25 Divide both eqns